# Mixture of Experts in Image Classification: What’s the Sweet Spot?

### v1

@@ Appendix B Routing analysis
Appendix C Detailed hyperparameters
Hyperparameter ViT (Touvron, Cord, and Jégou 2022) ConvNext (Liu et al. 2022)
@@ Appendix C Detailed hyperparameters
Table 11: ImageNet-1k and ImageNet-21k stochastic depth hyperparameter.
@@
Appendix D Impact of expert size and number
Figure 5: Exploring the interplay between the size and count of experts in a MoE layer for ConvNext-T on ImageNet-1K. Left, x-axis: number of experts. Y-axis: Accuracy on the given test-set. Size of points: MLP-ratio. Right, x-axis: MLP-ratio of experts, i.e., ratio between the number of hidden channels and the number of input/output channels. Y-axis: Accuracy on the given test-set. Size of points: number of experts. Baseline results (without MoE) are denoted by dotted lines. For this small dataset, we observe that MoE does not bring much improvement (see Figure 8 for bigger datasets), the best number of experts is 4 or 8, and the MoE can compensate for the performance loss associated with a smaller model size per sample.
Two critical tunable parameters within the MoE layers are the number of experts per layer and the size of each expert, where size corresponds to the mlp_ratio, representing the scaling factor of the hidden layer in the feed-forward network. Figure 5 shows the sensitivity of the results w.r.t. these parameters for ConvNext, with expert counts ranging from 4 to 16 and expert sizes from 1 to 4. Typically, this parameter is preset to 4, implying that using expert sizes less than 4 results in smaller per-sample layers compared to the original network. For comprehensive evaluation, we report the top-1 accuracy for three categories of datasets: independent and identically distributed (i.i.d.) (ImageNet, ImageNetV2 (Recht et al. 2019)), Out of Distribution (o.o.d.) (ImageNet-R (Hendrycks et al. 2021a), ImageNet-SK (Wang et al. 2019)), and adversarial (ImageNet-A (Hendrycks et al. 2021b)). The goal is to ascertain whether a configuration of many smaller experts or fewer larger experts yields superior performance.
Our observations on ConvNext, reveal that while MoE layers may negatively affect performance on o.o.d. datasets, they can maintain or slightly improve performance on i.i.d. datasets. Specifically, when the expert count reaches 16, performance deteriorates across all metrics. On the expert size aspect, smaller experts appear more sightly more resilient to increases in the number of experts. Our results indicate that an optimal configuration tends to involve 4 to 8 experts with a mlp_ratio of 2 or 4. Consequently, we adopted a configuration of 4 experts with a mlp_ratio of 2 for ImageNet-1K. For the ConvNext isotropic and ViT architecture, reducing the mlp_ratio always impairs results and even worst for ViT it produces training instabilities.
Appendix E On data-augmentation
As shown in the previous section, MoE is sensitive to the amount of data at hand. This sensitivity stems from the inherent structure of MoE, where increasing the number of experts results in each expert operating on a smaller data fraction. Consequently, each expert tends to learn specific data clusters, as discussed by Chen et al. (2022).
Modern training methodologies often employ robust data augmentation techniques to enhance model generalization. However, the interplay between these augmentations and the data clusters learned by the MoE model remains relatively uncharted. For instance, techniques like Mixup (Zhang et al. 2018), which blends images, could blur the distinctions between these clusters. Similarly, the Random Resize method, with its aggressive magnification capability, may distort the original data clusters. Understanding the precise influence of these augmentations on the MoE model’s learning process is essential, as it holds implications for the model’s ultimate performance. Some evidence of this can be seen in Table 13. While one training recipe might show MoE outperforming its dense counterparts (as indicated in the top rows of the table), a different, more effective recipe could lead to the opposite outcome (Bottom rows of the table).
Architecture # Params Per sample IN-1K
(×$\times 10^{6}$) #Paramsact Acc.
ViT-B (Touvron et al. 2021) 86.6 86.6 81.8
ViT-B-8 Every 2 Top 2 284.9 129.9 82.2
ViT-B (Touvron, Cord, and Jégou 2022) 86.6 86.6 82.8
ViT-B-8 Every 2 Top 2 284.9 129.9 82.5
Table 13: ViT-B trained on ImageNet-1k with different hyperparameters. In particular, for the top row (gray) we follow the training recipe from Deit (Touvron et al. 2021) and for the bottom one we follow Deit III (Touvron, Cord, and Jégou 2022).
Appendix F More model inspection
(a) 8 experts
(b) 16 experts
(c) 32 experts
Figure 7: Cumulated Distribution Function of the number of patches used per image for ConvNext-S trained on ImageNet-21k, with 8, 16, and 32 experts. Each colour corresponds to one expert. Sample detailed interpretation: for 8 experts, the “red” expert sees between 1 and 30 patches (out of 196) for 80% of the images on which it is active. In particular (and even more so when there are many experts), an expert frequently covers a limited part of the image. For instance, if the y-axis is close to 1 at x-axis=100 patches, covering more than 100 patches over 196 is rare: experts don’t specialize much. As the number of experts increases, this tendency of experts being used only on a few patches increases too. For 32 experts, most of them focus on less than 30 patches 80% of the time they are active.
Appendix G MoE vs Dense Table
@@
(×$\times 10^{6}$) #Paramsact Acc.
ViT-s/32(Riquelme et al. 2021) JFT-300M 36.5 36.5 73.73
ViT-s/32-32, Last 2(Riquelme et al. 2021) JFT-300M 166.7 $\approx 55$ 77.10
ViT-s/32-32, Every 2(Riquelme et al. 2021) JFT-300M 296.9 $\approx 70$ 77.10
ViT-S/16 (Touvron, Cord, and Jégou 2022) ImageNet-21k 22.0 22.0 82.6
@@
SwinV2-S (Hwang et al. 2022) ImageNet-21k 65.8 - 83.5
SwinV2-S-8 (Hwang et al. 2022) ImageNet-21k 173.3 65.8 84.5
SwinV2-S-16 (Hwang et al. 2022) ImageNet-21k 296.1 65.8 84.9
SwinV2-S-32 (Hwang et al. 2022) ImageNet-21k 296.1 65.8 84.7
ConvNeXt-S (Liu et al. 2022) ImageNet-21k 50.3 - 84.6
@@
ViT-B/16384absent384$\uparrow$ 384↑ 384 (Touvron, Cord, and Jégou 2022) ImageNet-21k 86.6 86.6 86.7
ViT-B/16 (Zhai et al. 2022) JFT-300M 86.6 86.6 84.9
ViT-B/16 (Riquelme et al. 2021) JFT-300M 100.5 100.5 84.15
V-MoE-B/16-32, Every 2 (Riquelme et al. 2021) JFT-300M 979.0 $\approx 200$ 85.3
V-MoE-B/16-32, Last 2 (Riquelme et al. 2021) JFT-300M 393.3 $\approx 110$ 85.4
SwinV2-B (Hwang et al. 2022) ImageNet-21k 109.3 109.3 85.1
SwinV2-B-8 (Hwang et al. 2022) ImageNet-21k 300.3 109.3 85.3
SwinV2-B-16 (Hwang et al. 2022) ImageNet-21k 518.7 109.3 85.5
SwinV2-B-32 (Hwang et al. 2022) ImageNet-21k 955.3 109.3 85.5
ConvNeXt-B (Liu et al. 2022) ImageNet-21k 88.6 88.6 85.8
@@
ViT-L/16 (Touvron, Cord, and Jégou 2022) ImageNet-21k 304.4 304.4 87.0
ViT-L/16 (Zhai et al. 2022) JFT-300M 323.1 323.1 87.7
ViT-L/16 (Riquelme et al. 2021) JFT-300M 323.1 323.1 87.1
V-MoE-L/16-32, Every 2 (Riquelme et al. 2021) JFT-300M 3446.0 $\approx 600$ 87.4
V-MoE-L/16-32, Last 2 (Riquelme et al. 2021) JFT-300M 843.6 $\approx 380$ 87.5
SoViT/14(Alabdulmohsin et al. 2023) JFT-3B 400 400 90.3
ViT-g/14(Zhai et al. 2022) JFT-3B 1B 1B 90.45
V-MoE/14 (Riquelme et al. 2021) JFT-3B 15B $\approx 1$B 90.35
Table 14: Experiments with pretraining on ImageNet-21k, JFT-300M or JFT-3B: performance with/without MoE. In the case of ConvNeXt-T, we get 0.6 improvement, while increasing the number of parameters but almost not the number of per-sample parameters. For ConvNeXt-S, there is a 0.3 improvement. There is an improvement for SwinV2 models, but these were starting lower. Overall, for strong/big models, there is no clear improvements. Results without citations correspond to our work. V-MoE results marked with a star () are results taken from their GitHub (tinyurl.com/yb8fze5u). See Figure 2 for an overview. For every model size examined, we present the accuracy metrics obtained after similar numbers of training iterations. Accuracy (Acc.) refers to Top-1 accuracy.
Figure 8: Top: Pareto-front for ImageNet21k, x-axis = number of activations per sample. ViT models have been presented in MoE versions only after additional pretraining, and are therefore not presented. MoE seems to be Pareto optimal for a number of activations per sample below 90M. Bottom: Pareto-front for ViT models on JFT-300M. Overall, MoE is never validated for a number of activations per sample above $\approx$ 100M.

### v2

@@ 5.6 Model Inspection
Expert Similarity: In order to quantify the similarity between experts, we use a method derived from the Jaccard index. For each image, we count the occurrences $c_{i}$ of each expert $E_{i}$. The similarity between two experts, $E_{i}$ and $E_{j}$, is defined as $S_{ij}=\tfrac{|E_{i}\cap E_{j}|}{|E_{i}\cup E_{j}|}=\tfrac{\sum^{E_{i}\cap E_{j}}c_{i}+c_{j}-|c_{i}-c_{j}|}{\sum^{E_{i}\cup E_{j}}c_{i}+c_{j}}$, where ${E_{i}\cap E_{j}}$ represents the images where both $E_{i}$ and $E_{j}$ appear together, while ${E_{i}\cup E_{j}}$ refers to all images where either $E_{i}$ or $E_{j}$ is present at least once. The absolute value term $|c_{i}-c_{j}|$ acts as a corrective factor to account for scenarios where one expert may be deployed for a minimal number of patches while the other is utilized for the majority, indicating a lack of synergy between the two experts. In the bottom panel of Fig.˜9, we observe a distinct pattern in the utilization of experts across the stages of the MoE. In the initial layers, there is a prevalent co-occurrence of multiple experts, which highlights a concentration on local image features across all the images. Conversely, as we analyze the network deeper, the frequency of expert co-occurrence diminishes. This shift indicates a transition in focus from local image elements to more semantically rich aspects, with individual experts concentrating on specific semantic attributes unique to certain images.
Are experts well partitioned and clustered?  For a general task such as ImageNet-1k, there is no clear, natural decomposition into experts, making it difficult to precisely assess if one expert decomposition is better than another, aside from comparing final accuracy. This comparison, however, provides limited insight into the quality of the experts. The literature (Chen et al., 2023; Mittal et al., 2022; Mohammed et al., 2022) suggests that, without explicit enforcement, MoE models struggle to separate tasks or find natural clusters beneath them. For instance, when a task separation was explicitly incorporated into the loss, there was a marked improvement in both accuracy and the ease with which experts could be pruned for specific tasks (Chen et al., 2023). Similarly, Mittal et al. (2022) found that the enforced decomposition resulted in substantial gains, while the absence of such enforcement left the model unable to discover optimal partitions. Consequently, we make three observations on ImageNet. First, as the number of experts increases, there is a corresponding rise in the number of experts actively involved. Even with 32 experts, almost half of the experts are used per image in the last MoE layers. This seems counterintuitive since ImageNet is centered around objects; at least for deep layers, it should plateau at some point. Instead, experts are not well aligned with classes and spread widely, as depicted in Fig.˜9. Second, an analysis of the cumulative distribution function (CDF) of experts’ engagement (Fig.˜8) reveals a prevalent trend in which most experts are assigned to less than 20 patches of an image, equating to less than 10% of the total image area, 50% of the time. Moreover, these patches are not necessarily contiguous. Third, when we look at the visualization of each expert, like in Fig.˜6, we observe a concentration of experts in the image for MoE layers near the output. The interpretation of each expert’s role becomes challenging: An expert does not seem to adhere closely to an object, or to a part of the location of an object.
Lack of Coherent Expert Partitioning Qualitative and quantitative analyses reveal no consistent expert partitioning in vision MoE, particularly as the number of experts increases. With few experts, routing weakly aligns with ImageNet classes, but as the scale grows, expert allocation becomes diffuse, many experts activate per image, each covering only a small portion of patches, and long-range consistency across class IDs becomes less clear. ConvNeXt shows some spatial biases, whereas ViT shows none; in both, assignments remain well load-balanced yet not semantically disentangled, suggesting that the observed gains stem from added capacity rather than interpretable specialization.
6 Conclusion
Implementing effective Mixture-of-Experts (MoE) models for image classification tasks remains a challenging and open problem. Our experiments show that, particularly in large-scale models, MoE for image classification offers limited gains over state-of-the-art methods in accuracy, robustness, generalization, or per-sample efficiency across both convolutional networks and vision transformers. This contrasts with NLP tasks, where MoE improves performance by enhancing capacity for knowledge storage and retrieval. In image classification, which relies more on feature extraction and extrapolation, MoE models show only modest benefits, mainly in smaller models with fewer activated parameters. These findings suggest that the effectiveness of MoE in image classification tasks is highly context- and scale-dependent, with advantages diminishing as model capacity grows.
Acknowledgements
The authors thank the reviewers for their insightful and constructive feedback, which has substantially improved the quality of this work. AL acknowledges support from the French National Research Agency (ANR) under Grant No. ANR-23-CPJ1-0099-01.
Appendix A Detailed Hyperparameters
We train all models with distributed data parallel (DDP) across devices and expert-parallelism within layers using Tutel (Hwang et al., 2023b) for efficient MoE implementation.
Hyperparameter ViT Touvron et al. (2022) ConvNeXt Liu et al. (2022)
@@ Appendix A Detailed Hyperparameters
Appendix B MoE vs Dense Table
Table 11: Experiments with pretraining on ImageNet-21k, JFT-300M or JFT-3B: performance with/without MoE. In the case of ConvNeXt-T, we get 0.6 improvement, while increasing the number of parameters, but almost not the number of per-sample parameters. For ConvNeXt-S, there is a 0.3 improvement. There is an improvement for SwinV2 models, but these were starting lower. Overall, for strong/big models, there is no clear improvement. Results without citations correspond to our work. V-MoE results marked with a star () are results taken from their GitHub (tinyurl.com/yb8fze5u). See Fig.˜4 for an overview. For every model size examined, we present the accuracy metrics obtained after similar numbers of training iterations. Accuracy (Acc.) refers to Top-1 accuracy.
@@ Appendix B MoE vs Dense Table
($\times 10^{6}$) #Paramsact Acc.
ViT-s/32(Riquelme et al., 2021) JFT-300M 36.5 36.5 73.73
ViT-s/32-32, Last 2(Riquelme et al., 2021) JFT-300M 166.7 $\approx 55$ 77.10
ViT-s/32-32, Every 2(Riquelme et al., 2021) JFT-300M 296.9 $\approx 70$ 77.10
ViT-S/16 (Touvron et al., 2022) ImageNet-21k 22.0 22.0 82.6
@@
SwinV2-S (Hwang et al., 2023a) ImageNet-21k 65.8 - 83.5
SwinV2-S-8 (Hwang et al., 2023a) ImageNet-21k 173.3 65.8 84.5
SwinV2-S-16 (Hwang et al., 2023a) ImageNet-21k 296.1 65.8 84.9
SwinV2-S-32 (Hwang et al., 2023a) ImageNet-21k 296.1 65.8 84.7
ConvNeXt-S (Liu et al., 2022) ImageNet-21k 50.3 - 84.6
@@
[…]
ViT-B/16 (Zhai et al., 2022) JFT-300M 86.6 86.6 84.9
ViT-B/16 (Riquelme et al., 2021) JFT-300M 100.5 100.5 84.15
V-MoE-B/16-32, Every 2 (Riquelme et al., 2021) JFT-300M 979.0 $\approx 200$ 85.3
V-MoE-B/16-32, Last 2 (Riquelme et al., 2021) JFT-300M 393.3 $\approx 110$ 85.4
SwinV2-B (Hwang et al., 2023a) ImageNet-21k 109.3 109.3 85.1
SwinV2-B-8 (Hwang et al., 2023a) ImageNet-21k 300.3 109.3 85.3
SwinV2-B-16 (Hwang et al., 2023a) ImageNet-21k 518.7 109.3 85.5
SwinV2-B-32 (Hwang et al., 2023a) ImageNet-21k 955.3 109.3 85.5
ConvNeXt-B (Liu et al., 2022) ImageNet-21k 88.6 88.6 85.8
@@
ViT-L/16 (Touvron et al., 2022) ImageNet-21k 304.4 304.4 87.0
ViT-L/16 (Zhai et al., 2022) JFT-300M 323.1 323.1 87.7
ViT-L/16 (Riquelme et al., 2021) JFT-300M 323.1 323.1 87.1
V-MoE-L/16-32, Every 2 (Riquelme et al., 2021) JFT-300M 3446.0 $\approx 600$ 87.4
V-MoE-L/16-32, Last 2 (Riquelme et al., 2021) JFT-300M 843.6 $\approx 380$ 87.5
SoViT/14(Alabdulmohsin et al., 2023) JFT-3B 400 400 90.3
ViT-g/14(Zhai et al., 2022) JFT-3B 1B 1B 90.45
V-MoE/14 (Riquelme et al., 2021) JFT-3B 15B $\approx 1$B 90.35
Appendix C Ade20k
Table 12: ImageNet-1K trained models
Backbone Input crop. #Params per sample Single scale
($\times 10^{6}$) #Params mIoU
ImageNet-1K trained models
ConvNeXt-T Liu et al. (2022) $512^{2}$ 28.6 - 46.0
ConvNeXt-T-4 Last 2 Top 1 $512^{2}$ 34.5 25.6 46.0
ConvNeXt-S Liu et al. (2022) $512^{2}$ 50 - 48.7
ConvNeXt-S-4 Last 2 Top 1 $512^{2}$ 56.1 47.3 48.38
ConvNeXt-B Liu et al. (2022) $512^{2}$ 88.6 - 49.1
ConvNeXt-B-4 Last 2 Top 1 $512^{2}$ 99.1 83.4 48.8
ViT-S Touvron et al. (2022) $512^{2}$ 88.6 - 45.8
ViT-S-8 Every 2 Top 2 $512^{2}$ 99.1 83.4 45.4
ViT-B Touvron et al. (2022) $512^{2}$ 88.6 - 49.0
ViT-B-8 Every 2 Top 2 $512^{2}$ 99.1 83.4 48.3
ImageNet-22K trained models
ConvNeXt-T Liu et al. (2022) $640^{2}$ 28.6 - 48.6
ConvNeXt-T-4 Last 2 Top 1 $640^{2}$ 34.5 25.6 48.9
ConvNeXt-S Liu et al. (2022) $640^{2}$ 50 - 51.0
ConvNeXt-S-4 Last 2 Top 1 $640^{2}$ 56.1 47.3 50.8
ConvNeXt-B Liu et al. (2022) $640^{2}$ 88.6 - 52.6
ConvNeXt-B-4 Last 2 Top 1 $640^{2}$ 99.1 83.4 52.0
ViT-S Touvron et al. (2022) $640^{2}$ 88.6 - 48.4
ViT-S-8 Every 2 Top 2 $640^{2}$ 99.1 83.4 48.7
ViT-B Touvron et al. (2022) $640^{2}$ 88.6 - 52.8
ViT-B-8 Every 2 Top 2 $640^{2}$ 99.1 83.4 52.4